OpenStudy (anonymous):

If the ellipse given by (x-3)^2/5^2 +(y-2)^2/3^2 is shifted two units up, what is the translated equation?

OpenStudy (anonymous):

\[\frac{ (x-3)^{2} }{ 5^{2} }+\frac{ (y-2)^{2} }{ 3^{2} }=1\]